AB INITIO ROTATION-VIBRATION SPECTRA OF X̃ ^{2}Σ^{+} MgNC

TINA ERICA ODAKA, TSUNEO HIRANO, Department of Chemistry, Faculty of Science, Ochanomizu University, 2-1-1 Otsuka, Bunkyo-ku, Tokyo 112-8610, Japan; \underline{PER JENSEN}^{a}, Institute for Astrophysics and Planetary Sciences, Ibaraki University, 2-1-1 Bunkyo, Mito 310-8512, Japan.

In 1986 Guélin et al.^{b} found, in radio astronomical observations, six transitions originating in the circumstellar envelope of the late-type carbon star IRC+10216. No assignment of these lines could be given at the time, but they were later identified^{c}^{,}^{d} as belonging to the rotational spectrum of the MgNC radical. Thus, MgNC became the first Mg-containing molecule to be identified in interstellar space. The only rotationally resolved, spectroscopic data presently available for X̃ ^{2}Σ^{+} MgNC comprise the rotational spectrum^{c,}^{e} together with a few vibronic bands, all originating in the vibronic ground state and belonging to the Ã ^{2}Π ← X̃ ^{2}Σ^{+} electronic transition.^{f} In the hope of stimulating further characterization of X̃ ^{2}Σ^{+} MgNC by high resolution spectroscopy, we report here ab initio simulations of its lowest X̃-state rotationvibration bands. The calculations are carried out with the MORBID program system,^{g} and they are based on a previously calculated potential energy function using ACPF method,^{h} supplemented with dipole moment surfaces computed with CASSCF/[TZ3P+f(Mg), aug-cc-pVQZ(N andC)].

aSpecial Visiting Professor of the Japanese Ministry of Education, Science and Culture, October-December 2000. Permanent address: FB 9 - Theoretische Chemie, Bergische Universität - Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany.

bM. Guélin, J. Cernicharo, C. Kahane, and J. Gomez-Gonzales, Astronomy & Astrophysics 157, L19 (1986)

cK. Kawaguchi, E. Kagi, T. Hirano, S. Takano, and S. Saito, Astrophys. J. 406, L39 (1993)

dK. Ishii, T. Hirano, U. Nagashima, B. Weis and K Yamashita, Astrophys. J. 410, L43 (1993)

eM. A. Anderson and L. M. Ziurys, Chem. Phys. Lett. 231, 164 (1994); E. Kagi, K. Kawaguchi, S. Takano, and T. Hirano, J. Chem. Phys. 104, 1263 (1996)

fR. R. Wright and T. A. Miller, J. Mol. Spectrosc. 194, 219 (1999)

gSee, for example, Section 15.4.7 of P. R. Bunker and P. Jensen, Molecular Symmetry and Spectroscopy, NRC Research Press, Ottawa, 1998, and references therein.

hT. Kinoshita and T. Hirano, 17th Austin Symposium on Molecular Structure, Abstracts, 48 (1998).